# Classical harmonic analysis viewed through the prism of noncommutative geometry

### v3

@@ 8 Examples of triples and Banach Fredholm modules
Suppose that $1<p<\infty$. The Banach odd Fredholm module over the algebra $\mathrm{C}_{0}(\mathbb{R})$ associated to the triple $(\mathrm{C}_{0}(\mathbb{R}),\mathrm{L}^{p}(\mathbb{R}),\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x})$ is $(\mathrm{L}^{p}(\mathbb{R}),\mathrm{i}\mathcal{H})$. For any function $f\in\mathrm{C}_{0}^{\infty}(\mathbb{R})$ satisfying (8.21), we have
(8.22) $\big\langle[f],(\mathrm{L}^{p}(\mathbb{R}),\pi,\mathrm{i}\mathcal{H})\big\rangle_{\mathrm{K}_{1}(\mathrm{C}_{0}(\mathbb{R})),\mathrm{K}^{1}(\mathrm{C}_{0}(\mathbb{R}),\mathscr{L}^{p})}=-\operatorname{wind}f,$
where is the winding number of $f$.
Proof : We let $F\overset{\mathrm{def}}{=}\mathrm{i}\mathcal{H}$. We have $F^{2}=\mathrm{Id}$. It is known [Uch78, Theorem 2 p. 17] that a function $f\in\mathrm{L}^{\infty}(\mathbb{R})$ belongs to the space $\mathrm{VMO}(\mathbb{R})$ of functions of vanishing mean oscillation if and only if the commutator $F,M_{f}]$ is a compact operator acting on the Banach space $\mathrm{L}^{p}(\mathbb{R})$. We infer that $(\mathrm{L}^{p}(\mathbb{R}),F)$ is a odd Fredholm module over the algebra $\mathrm{C}_{0}(\mathbb{R})$ of continuous functions that vanish at infinity, since $\mathrm{C}_{0}(\mathbb{R})$ is a subspace of $\mathrm{VMO}(\mathbb{R})$.
Moreover, observe that the map $P\overset{\mathrm{def}}{=}\frac{\mathrm{Id}+F}{2}$ identifies to the bounded projection $\mathrm{L}^{p}(\mathbb{R})\to\mathrm{L}^{p}(\mathbb{R})$ on the subspace $\mathrm{H}^{p}(\mathbb{R})$. Let $f\in\mathrm{C}_{0}(\mathbb{R})$ satisfying (8.21). This means that $f$ is an invertible element of the unitization $\mathrm{C}(X\cup\{\infty\})$ of the algebra $\mathrm{C}_{0}(X)$. The operator $PM_{f}P\colon P(\mathrm{L}^{p}(\mathbb{R}))\to P(\mathrm{L}^{p}(\mathbb{R}))$ of (4.1) (with $n=1$) identifies to the Toeplitz operator $T_{f}\colon\mathrm{H}^{p}(\mathbb{R})\to\mathrm{H}^{p}(\mathbb{R})$, $g\mapsto P(fg)$ with symbol $f$, where we identify the map $P$ with its corestriction $P|^{\mathrm{H}^{p}(\mathbb{R})}$. So the operator $T_{f}$ is Fredholm. Using Theorem 8.10, we deduce that
@@
Our approach is flexible. We can state variations with a possible weight, matricial versions and $\mathrm{UMD}$-vector-valued variants.
Remark 8.14
Suppose that $p=2$. If $0<q<\infty$, It is known that the commutator $F,M_{f}]$ is in the Schatten class $S^{q}(\mathrm{L}^{2}(\mathbb{R}))$ if and only if the function $f$ is in the Besov space $\mathrm{B}^{q}_{q,q}(\mathbb{R})$.
@@
acting on a subspace of the Banach space $\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})$. See [LaM89, p. 119]. The square of this operator is given by $D^{2}=\begin{bmatrix}-4\partial_{\overline{z}}\partial_{z}&0\\ 0&-4\partial_{z}\partial_{\overline{z}}\\ \end{bmatrix}=\begin{bmatrix}-\Delta&0\\ 0&-\Delta\\ \end{bmatrix}$. In the sequel, we use the notations $\partial\overset{\mathrm{def}}{=}2\partial_{z}$ and $\partial^{*}\overset{\mathrm{def}}{=}-2\partial_{\overline{z}}$. We start with a technical result.
@@
By [HvNVW18, Theorem 10.2.25 p. 391], note that the Laplacian $\Delta$ has a bounded $\mathrm{H}^{\infty}(\Sigma_{\theta})$ functional calculus for any angle $\theta>0$. Moreover, the Banach space $\mathrm{L}^{p}(\mathbb{C})$ is $\mathrm{UMD}$ by [HvNVW16, Proposition 4.2.15 p. 291], hence has the triangular contraction property $(\Delta)$ by [HvNVW18, Theorem 7.5.9 p. 137]. We deduce by [HvNVW18, Theorem 10.3.4 (2) p. 401] that the operator $\Delta$ is $R$-sectorial. By [HvNVW18, Example 10.3.5 p. 402] applied with $\alpha=\frac{1}{2}$ and $\beta=1$, we infer that the set
@@
Suppose that $1<p<\infty$. The unbounded operator $D$ admits a bounded $\mathrm{H}^{\infty}(\Sigma_{\theta}^{\mathrm{bi}})$ functional calculus on the Banach space $\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})$ for some $0<\theta<\frac{\pi}{2}$.

### v4

@@ 8 Examples of triples and Banach Fredholm modules
Suppose that $1<p<\infty$. The triple $(\mathrm{L}^{p}(\mathbb{R}),\pi,\mathrm{i}\mathcal{H})$ is an odd Banach Fredholm module over the algebra $\mathrm{C}_{0}(\mathbb{R})$. For any function $f\in\mathrm{C}_{0}^{\infty}(\mathbb{R})$ satisfying (8.34), we have
(8.35) $\big\langle[f],(\mathrm{L}^{p}(\mathbb{R}),\pi,\mathrm{i}\mathcal{H})\big\rangle_{\mathrm{K}_{1}(\mathrm{C}_{0}(\mathbb{R})),\mathrm{K}^{1}(\mathrm{C}_{0}(\mathbb{R}),\mathscr{L}^{p})}=-\operatorname{wind}f.$
Proof : We let $F\overset{\mathrm{def}}{=}\mathrm{i}\mathcal{H}$. We have $F^{2}=\mathrm{Id}$. It is known [Uch78, Theorem 2 p. 17] that a function $f\in\mathrm{L}^{\infty}(\mathbb{R})$ belongs to the space $\mathrm{VMO}(\mathbb{R})$ of functions of vanishing mean oscillation if and only if the commutator $F,M_{f}]$ is a compact operator acting on the Banach space $\mathrm{L}^{p}(\mathbb{R})$. Since $\mathrm{C}_{0}(\mathbb{R})$ is a subspace of $\mathrm{VMO}(\mathbb{R})$ by [Ste93, 6.8 (a) p. 180], we infer that $(\mathrm{L}^{p}(\mathbb{R}),\pi,F)$ is a odd Fredholm module over the algebra $\mathrm{C}_{0}(\mathbb{R})$ of continuous functions that vanish at infinity.
Let $f\in\mathrm{C}_{0}(\mathbb{R})$ satisfying (8.34). This means that $f$ is an invertible element of the unitization $\mathrm{C}(X\cup\{\infty\})$ of the algebra $\mathrm{C}_{0}(X)$. The operator $PM_{f}P\colon P(\mathrm{L}^{p}(\mathbb{R}))\to P(\mathrm{L}^{p}(\mathbb{R}))$ of (4.4) (with $n=1$) identifies to the Toeplitz operator $T_{f}\colon\mathrm{H}^{p}(\mathbb{R})\to\mathrm{H}^{p}(\mathbb{R})$, $g\mapsto P(fg)$ with symbol $f$, where we identify the map $P$ with its corestriction $P|^{\mathrm{H}^{p}(\mathbb{R})}$. So the operator $T_{f}$ is Fredholm. Using Theorem 8.19, we deduce that
@@
Our approach is flexible. We can state variations with a possible weight [HMW73], matricial versions and $\mathrm{UMD}$-vector-valued variants.
@@
The following result is an analogue of Proposition 8.8.
Proposition 8.23
Suppose that $1<p<\infty$. Consider the unbounded operator $D\overset{\mathrm{def}}{=}\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ on $\mathrm{L}^{p}(\mathbb{R})$. We have $|D|=\mathcal{H}\,\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$. Moreover, for any $f\in\mathrm{C}^{\infty}_{c}(\mathbb{R})$, the commutator $[|D|,f]$ induces a bounded operator on $\mathrm{L}^{p}(\mathbb{R})$ and
(8.36) $\big([|D|,f]g\big)(x)=\frac{1}{\pi}\operatorname{p.v.}\int_{\mathbb{R}}\frac{f(x)-f(y)}{(x-y)^{2}}\,g(y)\,\mathrm{d}y,\qquad g\in\mathcal{S}(\mathbb{R}).$
Finally, we have
(8.37) $\big\|[|D|,f]\big\|_{\mathrm{L}^{p}(\mathbb{R})\to\mathrm{L}^{p}(\mathbb{R})}\lesssim_{p}\|f^{\prime}\|_{\mathrm{L}^{\infty}(\mathbb{R})}.$
Proof : First proof. The operator $|D|$ is the unbounded Fourier multiplier defined by the symbol $|\xi|$. The first assertion is a consequence of (8.32). Fix $f\in\mathrm{C}^{\infty}_{c}(\mathbb{R})$ and $g\in\mathcal{S}(\mathbb{R})$. Using $|D|=\mathcal{H}\,\frac{\mathrm{d}}{\mathrm{d}x}$,
$|D|,M_{f}]g=\mathcal{H}\,(fg)^{\prime}-f\,\mathcal{H}\,g^{\prime}=\mathcal{H}(f^{\prime}g)+\big(\mathcal{H}(fg^{\prime})-f\,\mathcal{H}g^{\prime}\big)=\mathcal{H}(f^{\prime}g)+[\mathcal{H},M_{f}]g^{\prime}.$
Write both terms using the principal value representation of $\mathcal{H}$ and integrate by parts in the second one. More precisely, we have
[…]
$=\frac{1}{\pi}\operatorname{p.v.}\int_{\mathbb{R}}\frac{f(y)-f(x)}{x-y}g^{\prime}(y)\mathop{}\mathopen{}\mathrm{d}y=-\frac{1}{\pi}\operatorname{p.v.}\int_{\mathbb{R}}\Big(\frac{f^{\prime}(y)}{x-y}+\frac{f(y)-f(x)}{(x-y)^{2}}\Big)g(y)\mathop{}\mathopen{}\mathrm{d}y,$
[…]
(8.38) $K_{f}(x,y)\overset{\mathrm{def}}{=}\frac{1}{\pi}\frac{f(x)-f(y)}{(x-y)^{2}},\quad x\neq y.$
By the mean value theorem, we see that
[…]
Differentiating with respect to $x$ and using $\partial_{x}(x-y)^{-2}=-2(x-y)^{-3}$, we get
$\partial_{x}K_{f}(x,y)=\frac{1}{\pi}\left(\frac{f^{\prime}(x)}{(x-y)^{2}}-2\,\frac{f(x)-f(y)}{(x-y)^{3}}\right).$
Similarly, since $\partial_{y}(x-y)^{-2}=2(x-y)^{-3}$, we have
$\partial_{y}K_{f}(x,y)=\frac{1}{\pi}\left(-\frac{f^{\prime}(y)}{(x-y)^{2}}+2\,\frac{f(x)-f(y)}{(x-y)^{3}}\right).$
By the mean value theorem, we see that
$\left|2\,\frac{f(x)-f(y)}{(x-y)^{3}}\right|\leqslant 2\frac{\left\|f^{\prime}\right\|_{\mathrm{L}^{\infty}(\mathbb{R})}|x-y|}{|x-y|^{3}}=\frac{2\left\|f^{\prime}\right\|_{\mathrm{L}^{\infty}(\mathbb{R})}}{|x-y|^{2}}.$
It follows that
$|\partial_{x}K_{f}(x,y)|\leq\frac{1}{\pi}\left(\frac{\left\|f^{\prime}\right\|_{\mathrm{L}^{\infty}(\mathbb{R})}}{|x-y|^{2}}+\frac{2\left\|f^{\prime}\right\|_{\mathrm{L}^{\infty}(\mathbb{R})}}{|x-y|^{2}}\right)\lesssim\frac{\left\|f^{\prime}\right\|_{\mathrm{L}^{\infty}(\mathbb{R})}}{|x-y|^{2}},$
and similarly
$|\partial_{y}K_{f}(x,y)|\lesssim\frac{\left\|f^{\prime}\right\|_{\mathrm{L}^{\infty}(\mathbb{R})}}{|x-y|^{2}}.$
Adding the two bounds yields
$|\partial_{x}K_{f}(x,y)|+|\partial_{y}K_{f}(x,y)|\lesssim\frac{\left\|f^{\prime}\right\|_{\mathrm{L}^{\infty}(\mathbb{R})}}{|x-y|^{2}}.$
Hence $K_{f}$ is a Calderón–Zygmund kernel with constants $\lesssim\|f^{\prime}\|_{\mathrm{L}^{\infty}}$, so the principal value operator (8.36) extends boundedly to $\mathrm{L}^{p}(\mathbb{R})$ with the estimate (8.37) by the classical Calderón–Zygmund theory.
Second proof of the first assertion. Observe that $|D|$ is a pseudo-differential operator of order $1$. Moreover, recall that if $f\in\mathrm{C}_{c}^{\infty}(\mathbb{R})$ then $M_{f}$ is a differential operator of order 0, hence a pseudo-differential operator of order 0 by [Dieu88, Remark 23.26.5]. Consequently, according to [ADV25, Theorem H.9 p. 241] the commutator $|D|,M_{f}]$ is a pseudo-differential operator of order 0. Such operator induces a bounded operator on the Banach space $\mathrm{L}^{p}(\mathbb{R})$ by [RuT10, Theorem 2.6.22 p. 296] or [ADV25, Theorem H.10 p. 242].
@@
acting on the subspace $\mathrm{W}^{1,p}(\mathbb{C})\oplus\mathrm{W}^{1,p}(\mathbb{C})$ of the Banach space $\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})$. The square of this operator is given by $D^{2}=\begin{bmatrix}-4\partial_{\overline{z}}\partial_{z}&0\\ 0&-4\partial_{z}\partial_{\overline{z}}\\ \end{bmatrix}=\begin{bmatrix}-\Delta&0\\ 0&-\Delta\\ \end{bmatrix}$. In the sequel, we use the notations $\partial\overset{\mathrm{def}}{=}2\partial_{z}$ and $\partial^{*}\overset{\mathrm{def}}{=}-2\partial_{\overline{z}}$. See [LaM89, p. 119] for more information on this operator. We can introduce the isometric homomorphism $\pi\colon\mathrm{L}^{\infty}(\mathbb{C})\to\mathrm{B}(\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C}))$, $f\mapsto\begin{bmatrix}M_{f}&0\\ 0&M_{f}\\ \end{bmatrix}$.
We begin by determining the Lipschitz algebra.
Proposition 8.24
Suppose that $1<p<\infty$. Consider the operator $D\overset{\mathrm{def}}{=}\begin{bmatrix}0&-2\partial_{\overline{z}}\\ 2\partial_{z}&0\end{bmatrix}$ with domain $\operatorname{dom}D=\mathrm{W}^{1,p}(\mathbb{C})\oplus\mathrm{W}^{1,p}(\mathbb{C})$. Then
(8.40) $\operatorname{\mathrm{Lip}}_{D}(\mathrm{L}^{\infty}(\mathbb{C}))=\mathrm{W}^{1,\infty}(\mathbb{C}).$
Moreover, for any function $f\in\mathrm{W}^{1,\infty}(\mathbb{C})$ we have
(8.41) $D,f]=\begin{bmatrix}0&-2M_{\partial_{\overline{z}}f}\\ 2M_{\partial_{z}f}&0\end{bmatrix}$
and
(8.42) $\big\|[D,\pi(f)]\big\|_{\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})\to\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})}=2\max\Big\{\left\|\partial_{z}f\right\|_{\mathrm{L}^{\infty}(\mathbb{C})},\left\|\partial_{\overline{z}}f\right\|_{\mathrm{L}^{\infty}(\mathbb{C})}\Big\}.$
Proof : If $f\in\mathrm{W}^{1,\infty}(\mathbb{C})$, then by then $M_{f}(\mathrm{W}^{1,p}(\mathbb{C}))\subset\mathrm{W}^{1,p}(\mathbb{C})$. Hence $\operatorname{dom}D$ is invariant under the operator $\pi(f)$. For any $u,v\in\mathrm{C}_{c}^{\infty}(\mathbb{C})$, we have
(8.43) $\big[D,f\big]=\begin{bmatrix}0&-2\partial_{\overline{z}}\\ 2\partial_{z}&0\\ \end{bmatrix}\begin{bmatrix}M_{f}&0\\ 0&M_{f}\\ \end{bmatrix}-\begin{bmatrix}M_{f}&0\\ 0&M_{f}\\ \end{bmatrix}\begin{bmatrix}0&-2\partial_{\overline{z}}\\ 2\partial_{z}&0\\ \end{bmatrix}$
$=\begin{bmatrix}0&-2\partial_{\overline{z}}M_{f}\\ 2\partial_{z}M_{f}&0\\ \end{bmatrix}+\begin{bmatrix}0&2M_{f}\partial_{\overline{z}}\\ -2M_{f}\partial_{z}&0\\ \end{bmatrix}=\begin{bmatrix}0&-2[\partial_{\overline{z}},M_{f}]\\ 2[\partial_{z},M_{f}]&0\end{bmatrix}$.
Consequently, we have
[…]
Hence, we obtain (8.41). Since the functions $\partial_{\overline{z}}f$ and $\partial_{z}f$ belong to $\mathrm{L}^{\infty}(\mathbb{C})$, we infer that this operator admits a bounded extension. We deduce that $\mathrm{W}^{1,\infty}(\mathbb{C})\subset\operatorname{\mathrm{Lip}}_{D}(\mathrm{L}^{\infty}(\mathbb{C}))$. The inequality $\big\|[D,f]\big\|\leqslant 2\max\Big\{\left\|\partial_{z}f\right\|_{\mathrm{L}^{\infty}(\mathbb{C})},\left\|\partial_{\overline{z}}f\right\|_{\mathrm{L}^{\infty}(\mathbb{C})}\Big\}$ follows from a simple computation:
$\left\|\big(-2(\partial_{\overline{z}}f)v,2(\partial_{z}f)u\big)\right\|_{\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})}=\big(\left\|2(\partial_{\overline{z}}f)v\right\|_{\mathrm{L}^{p}(\mathbb{C})}^{p}+\left\|2(\partial_{z}f)u\right\|_{\mathrm{L}^{p}(\mathbb{C})}^{p}\big)^{\frac{1}{p}}$
$\leqslant 2\max\Big\{\left\|\partial_{z}f\right\|_{\mathrm{L}^{\infty}(\mathbb{C})},\left\|\partial_{\overline{z}}f\right\|_{\mathrm{L}^{\infty}(\mathbb{C})}\Big\}\big(\left\|v\right\|_{\mathrm{L}^{p}(\mathbb{C})}^{p}+\left\|u\right\|_{\mathrm{L}^{p}(\mathbb{C})}^{p}\big)^{\frac{1}{p}}$
$=2\max\Big\{\left\|\partial_{z}f\right\|_{\mathrm{L}^{\infty}(\mathbb{C})},\left\|\partial_{\overline{z}}f\right\|_{\mathrm{L}^{\infty}(\mathbb{C})}\Big\}\left\|(u,v)\right\|_{\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})}.$
By considering $(u,0)$ or $(0,u)$, we obtain $\big\|[D,f]\big\|\geqslant 2\max\Big\{\left\|\partial_{z}f\right\|_{\mathrm{L}^{\infty}(\mathbb{C})},\left\|\partial_{\overline{z}}f\right\|_{\mathrm{L}^{\infty}(\mathbb{C})}\Big\}$.
Now, suppose that $f\in\operatorname{\mathrm{Lip}}_{D}(\mathrm{L}^{\infty}(\mathbb{C}))$. By definition, the operator $\pi(f)$ leaves $\operatorname{dom}D=\mathrm{W}^{1,p}(\mathbb{C})\oplus\mathrm{W}^{1,p}(\mathbb{C})$ invariant and $D$ induces a bounded operator on $\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})$. Writing
$D,f]=\begin{bmatrix}0&-2[\partial_{\overline{z}},M_{f}]\\ 2[\partial_{z},M_{f}]&0\end{bmatrix}$,
we see that $\partial_{z},M_{f}]$ and $\partial_{\overline{z}},M_{f}]$ are bounded on $\mathrm{L}^{p}(\mathbb{C})$. Since $\partial_{x}=\partial_{z}+\partial_{\overline{z}}$ and $\partial_{y}=\tfrac{1}{\mathrm{i}}\,(\partial_{z}-\partial_{\overline{z}})$, it follows that $\partial_{x},M_{f}]$ and $\partial_{y},M_{f}]$ are bounded on $\mathrm{L}^{p}(\mathbb{C})$. Fix $j\in\{1,2\}$ and apply Proposition 8.2 with $R=M_{f}$ and the one-parameter translation group $(\tau_{te_{j}})_{t\in\mathbb{R}}$ in direction $e_{j}$ with generator $\partial_{j}$. We obtain, for all small $t\neq 0$,
[…]
Consequently, we have
$\big\|\Delta_{t}^{(j)}f\big\|_{\mathrm{L}^{\infty}(\mathbb{C})}=\big\|M_{\Delta_{t}^{(j)}f}\big\|_{\mathrm{L}^{p}(\mathbb{C})\to\mathrm{L}^{p}(\mathbb{C})}=\left\|\int_{0}^{1}\tau_{ste_{j}}[\partial_{x_{j}},M_{f}]\,\tau_{-ste_{j}}\mathop{}\mathopen{}\mathrm{d}s\right\|_{\mathrm{L}^{p}(\mathbb{C})\to\mathrm{L}^{p}(\mathbb{C})}$
[…]
By Banach-Alaoglu theorem, there exist $t_{\alpha}\to 0$ and $g_{j}\in\mathrm{L}^{\infty}(\mathbb{C})$ such that $\Delta_{t_{\alpha}}^{(j)}f\overset{}{\longrightarrow}g_{j}$ in the weak* topology of the dual space $\mathrm{L}^{\infty}(\mathbb{C})$. For any $\varphi\in\mathrm{C}_{c}^{\infty}(\mathbb{C})$, we have
$\int_{\mathbb{C}}(\Delta_{t}^{(j)}f)\varphi=\int_{\mathbb{C}}f\frac{\varphi(\cdot-te_{j})-\varphi}{t}\xrightarrow[t\to 0]{}-\int_{\mathbb{C}}f(\partial_{x_{j}}\varphi),$
since $\frac{\varphi(\cdot-te_{j})-\varphi}{t}\to-\partial_{x_{j}}\varphi$ in $\mathrm{L}^{1}(\mathbb{R})$ as $t\to 0$. Passing to the weak* limit gives $\int g_{j}\varphi=-\int f\,\partial_{x_{j}}\varphi$. Thus $g_{j}=\partial_{x_{j}}f$ in $\mathcal{D}^{\prime}$. Consequently $\partial_{x_{j}}f\in\mathrm{L}^{\infty}(\mathbb{C})$. We conclude that $f\in\mathrm{W}^{1,\infty}(\mathbb{C})$.
@@
By Example 2.5, the Laplacian $\Delta$ admits a bounded $\mathrm{H}^{\infty}(\Sigma_{\theta})$ functional calculus for any angle $\theta>0$. Moreover, the Banach space $\mathrm{L}^{p}(\mathbb{C})$ is $\mathrm{UMD}$ by [HvNVW16, Proposition 4.2.15 p. 291], hence has the triangular contraction property $(\Delta)$ by [HvNVW18, Theorem 7.5.9 p. 137]. We deduce by [HvNVW18, Theorem 10.3.4 (2) p. 401] that the operator $\Delta$ is $R$-sectorial. By [HvNVW18, Example 10.3.5 p. 402] applied with $\alpha=\frac{1}{2}$ and $\beta=1$, we infer that the set
@@
Suppose that $1<p<\infty$. The unbounded operator $D$ admits a bounded $\mathrm{H}^{\infty}(\Sigma_{\theta}^{\mathrm{bi}})$ functional calculus on the Banach space $\mathrm{L}^{p}(\mathbb{C})\oplus\mathrm{L}^{p}(\mathbb{C})$ for some angle $\theta\in(0,\frac{\pi}{2})$.